OpenStudy (anonymous):

slope is -2 and points are (4,6)

OpenStudy (anonymous):

Yes, what about them?

OpenStudy (anonymous):

how do i write the equation

OpenStudy (anonymous):

What equation?

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

Write the equation of the line that passes through the y-axis at -2, and has points of (4,6)

OpenStudy (anonymous):

y-y1=m(x-x1) so use this and form it into y=mx+b

OpenStudy (anonymous):

y=mx+b yees

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

6=-2(4)+(-2)

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

what next?

OpenStudy (anonymous):

idk I think I might have done that wrong idk

OpenStudy (anonymous):

...

OpenStudy (anonymous):

Yeah lol I did that wrong

OpenStudy (anonymous):

Why do people take away medals, I guess you need to give the answer to get an actual one, which isn't the point of the site >_>.

OpenStudy (anonymous):

Sorry I'm a stereotypical blonde

OpenStudy (anonymous):

@iambatman I gave you a medal bud(:

OpenStudy (anonymous):

m= slope Points given = x1, y1 Rest of the variables you leave alone x, y. Thank you @Kenzieloo :)